# Variational Continual Test-Time Adaptation

### v1

@@ Abstract
The prior drift is crucial in Continual Test-Time Adaptation (CTTA) methods that only use unlabeled test data, as it can cause significant error propagation. In this paper, we introduce VCoTTA, a variational Bayesian approach to measure uncertainties in CTTA. At the source stage, we transform a pre-trained deterministic model into a Bayesian Neural Network (BNN) via a variational warm-up strategy, injecting uncertainties into the model. During the testing time, we employ a mean-teacher update strategy using variational inference for the student model and exponential moving average for the teacher model. Our novel approach updates the student model by combining priors from both the source and teacher models. The evidence lower bound is formulated as the cross-entropy between the student and teacher models, along with the Kullback-Leibler (KL) divergence of the prior mixture. Experimental results on three datasets demonstrate the method’s effectiveness in mitigating prior drift within the CTTA framework.
Machine Learning, ICML
1 Introduction
Continual Test-Time Adaptation (CTTA) (Wang et al., 2022) aims to enable a model to accommodate a sequence of distinct distribution shifts during the testing time (See Fig. 1(a)), making it applicable to various risk-sensitive applications in open environments, such as autonomous driving and medical imaging. However, real-world non-stationary test data exhibit high uncertainty in their temporal dynamics (Huang et al., 2022), presenting challenges related to error accumulation (Wang et al., 2022). Previous CTTA studies rely on methods that enforce prediction confidence, such as entropy minimization. However, these approaches often lead to predictions that are overly confident and less well-calibrated, thus limiting the model’s ability to quantify risks during predictions. The reliable estimation of uncertainty becomes particularly crucial in the context of continual distribution shift (Ovadia et al., 2019). Therefore, it is meaningful to design a model capable of encoding the uncertainty associated with temporal dynamics and effectively handling distribution shifts. The objective of this paper is to devise a test-time adaptation procedure that not only enhances predictive accuracy under distribution shifts but also provides reliable uncertainty estimates.
(a) CTTA task
(b) Unreliable prior in CTTA
Figure 1: A CTTA model is updated using pseudo labels, which may result in error accumulations and make the prior unreliable.
@@ 1 Introduction
To address the above problem, we refer to the Bayesian Inference (BI) approach (Box & Tiao, 2011), which retains a distribution over model parameters that indicates the plausibility of different settings given the observed data, and it has been witnessed as effective in traditional continual learning tasks (Nguyen et al., 2018). In Bayesian continual learning, the posterior in the last learning task is set to be the current prior which will be multiplied by the current likelihood. This kind of prior transmission is designed to reduce catastrophic forgetting in continual learning (Liu et al., ; Sun et al., 2022; Du et al., 2023; Lyu et al., 2023, 2021). However, this is not feasible in CTTA because unlabeled data may introduce significant prior drift. That is, as shown in Fig. 1(b), an unreliable prior may lead to a poor posterior, which may then propagate errors to the next inference, leading to the accumulation of errors.
@@
Thus, we delve into mitigating prior drift in CTTA within a BI framework in this paper. To approximate the intractable likelihood in Bayesian inference, we adopt to use online Variational Inference (VI) (Wang et al., 2011; Sato, 2001), and accordingly name our method Variational Continual Test-Time Adaptation (VCoTTA). At the source stage, we first transform a pretrained deterministic model into a Bayesian Neural Network (BNN) by a variational warm-up strategy, where the local reparameterization trick (Kingma et al., 2015) is used to inject uncertainties into the source model. During the testing phase, we employ a mean-teacher update strategy, where the student model is updated via VI and the teacher model is updated by the exponential moving average. Specifically, for the update of the student model, we propose to use a mixture of priors from both the source and teacher models, then the Evidence Lower BOund (ELBO) becomes the cross-entropy between the student and teachers plus the KL divergence of the prior mixture. We demonstrate the effectiveness of the proposed method on three datasets, and the results show that the proposed method can mitigate the prior drift in CTTA and obtain clear performance improvements.
@@
  1. 1.
    This paper develops VCoTTA, a simple yet general framework for continual test-time adaptation that leverages online VI within BNN.
  2. 2.
    We propose to transform an off-the-shelf model into a BNN via a variational warm-up strategy, which injects uncertainties into the model.
  3. 3.
    We build a mean-teacher structure for CTTA, and propose a strategy to blend the teacher’s prior with the source’s prior to mitigate prior drift.
2 Related Work
2.1 Continual Test-Time Adaptation
Test-Time Adaptation (TTA) enables the model to dynamically adjust to the characteristics of the test data, i.e. target domain, in a source-free and online manner (Jain & Learned-Miller, 2011; Sun et al., 2020; Wang et al., 2020). Previous works have enhanced TTA performance through the designs of unsupervised loss (Mummadi et al., 2021; Zhang et al., 2022; Liu et al., 2021; Choi et al., 2022; Chen et al., 2022; Gandelsman et al., 2022). These endeavours primarily focus on enhancing adaptation within a fixed target domain, representing a single-domain TTA setup. In this scenario, the model adapts to a specific target domain and then resets to its original pre-trained state with the source domain, prepared for the next target domain adaptation.
@@ 2.1 Continual Test-Time Adaptation
Recently, CTTA (Wang et al., 2022) has been introduced to tackle TTA within a continuously changing target domain, involving long-term adaptation. This configuration often grapples with the challenge of error accumulation (Tarvainen & Valpola, 2017; Wang et al., 2022). Specifically, prolonged exposure to unsupervised loss from unlabeled test data during long-term adaptation may result in significant error accumulation. Additionally, as the model is intent on learning new knowledge, it is prone to forgetting source knowledge, which poses challenges when accurately classifying test samples similar to the source distribution.
To solve the two challenges, the majority of the existing methods focus on improving the confidence of the source model during the testing phase. These methods employ the mean-teacher architecture (Tarvainen & Valpola, 2017) to mitigate error accumulation, where the student learns to align with the teacher and the teacher updates via moving average with the student. As to the challenge of forgetting source knowledge, some methods adopt augmentation-averaged predictions (Wang et al., 2022; Brahma & Rai, 2023; Döbler et al., 2023; Yang et al., 2023) for the teacher model, strengthening the teacher’s confidence to reduce the influence from highly out-of-distribution samples. Some methods, such as Döbler et al. (2023); Chakrabarty et al. (2023), propose to adopt the contrastive loss to maintain the already learnt semantic information. Some methods believe that the source model is more reliable, thus they are designed to restore the source parameters (Wang et al., 2022; Brahma & Rai, 2023). Though the above methods keep the model from confusion of vague pseudo labels, they may suffer from overly confident predictions that are less calibrated. To mitigate this issue, it is helpful to estimate the uncertainty in the neural network.
2.2 Bayesian Neural Network
@@
Bayesian framework is natural to incorporate past knowledge and sequentially update the belief with new data (Zhao et al., 2022). The bulk of work on Bayesian deep learning has focused on scalable approximate inference methods. These methods include stochastic VI (Hernández-Lobato & Adams, 2015; Louizos & Welling, 2017), dropout (Gal & Ghahramani, 2016; Kingma et al., 2015) and Laplace approximation (Ritter et al., 2018; Friston et al., 2007) etc., and leveraging the stochastic gradient descent (SGD) trajectory, either for a deterministic approximation or sampling. In a BNN, we specify a prior $p(\boldsymbol{\theta})$ over the neural network parameters, and compute the posterior distribution over parameters conditioned on training data, $p(\boldsymbol{\theta}|\mathcal{D})\propto p(\boldsymbol{\theta})p(\mathcal{D}|\boldsymbol{\theta})$. This procedure should give considerable advantages for reasoning about predictive uncertainty, which is especially relevant in the small-data setting.
@@
Crucially, when performing Bayesian inference, we need to choose a prior distribution that accurately reflects the prior beliefs about the model parameters before seeing any data (Gelman et al., 1995; Fortuin et al., 2021). In conventional static machine learning, the most common choice for the prior distribution over the BNN weights is the simplest one: the isotropic Gaussian distribution. However, this choice has been proved indeed suboptimal for BNNs (Fortuin et al., 2021). Recently, some studies estimate uncertainty in continual learning within a BNN framework, such as Nguyen et al. (2018); Ebrahimi et al. (2019); Farquhar & Gal (2019); Kurle et al. (2019). They set the current prior to the previous posterior to mitigate catastrophic forgetting. However, the prior transmission is not reliable in the unsupervised CTTA task. Any prior mistakes will be enlarged by adaptation progress, manifesting error accumulation, which can be seen in Fig. 2. To solve the prior drift problem, in this paper, we propose a prior mixture method based on VI.
3 Variational CTTA
@@
We start from the BI in typical continual learning, where the model aims to learn multiple classification tasks in sequence. Let $\mathcal{D}=\{(x_{n},y_{n})\}_{n=1}^{N}$ be the training set, where $x_{n}$ and $y_{n}$ denotes the training sample and the corresponding class label. The task $t$ is to learn a direct posterior approximation over the model parameter $\boldsymbol{\theta}$ as follows.
@@
where $p(\boldsymbol{\theta}|\mathcal{D}_{1:t})$ denotes the posterior of sequential tasks on the learned parameter and $p(\mathcal{D}_{t}|\boldsymbol{\theta})$ is the likelihood of the current task. The current prior $p_{t}(\boldsymbol{\theta})$ is regarded as the given knowledge. Nguyen et al. (2018) propose that this current prior can be the posterior learned in the last task, i.e., $p_{t}(\boldsymbol{\theta})=p(\boldsymbol{\theta}|\mathcal{D}_{1:t-1})$, thus the inference becomes
@@
This process can be shown in Fig. 2(a).
In contrast to continual learning, CTTA faces a sequence of learning tasks in test time without any label information, requiring the model to adapt to each novel domain sequentially. The CTTA model is first trained on a source dataset $\mathcal{D}_{0}$, and then adapted to unlabeled test domains starting from $\mathcal{U}_{1}$. For the $t$-th adaptation, we have
@@
Similarly, we can set the last posterior to be the current prior, i.e., $p_{t}(\boldsymbol{\theta})=p(\boldsymbol{\theta}|\mathcal{U}_{1:t-1}\cup\mathcal{D}_{0})$ and $p_{1}(\boldsymbol{\theta})=p(\boldsymbol{\theta}|\mathcal{D}_{0})$. However, the adaption with the unlabeled test data will cause significant error accumulation. In other words, an unreliable prior $p_{t}(\boldsymbol{\theta})$ will make the current posterior even less trustworthy. Moreover, the joint likelihood $p(\mathcal{U}_{t}|\boldsymbol{\theta})$ for $t>0$ is intractable on unlabeled data.
To make Bayesian inference feasible, in this paper, we transform the question to an easy-to-compute form. Referring to Grandvalet & Bengio (2004), the unsupervised inference can be transformed into
@@
where $H$ denotes the conditional entropy and $\lambda$ is a scalar hyperparameter to weigh the entropy term. This simple form reveals that the prior belief about the conditional entropy of labels is given by the inputs. The observation of unlabeled test inputs $\mathcal{U}$ provides information on the drift of the input distribution, which can be used to update the belief over the learned parameters $\boldsymbol{\theta}$ through Eq. (4). Consequently, this allows the utilization of unlabeled data for Bayesian inference. More detailed derivations can be seen in Appendix A.
(a) BI in continual learning
(b) BI in CTTA
Figure 2: Bayesian inference comparison between continual learning and CTTA. We find the traditional prior transmission is infeasible in CTTA because of the prior shift from unlabeled data. In our method, we place CTTA in a mean-teacher structure, and design BI in CTTA using a mixture of teacher prior and source prior. The next teacher prior is updated by the exponential moving average.
In a BNN, the posterior distribution is often intractable and some approximation methods are required, even when calculating the initial posterior. In this paper, we leverage online VI, as it typically outperforms the other methods for complex models in the static setting (Bui et al., 2016). VI defines a variational distribution $q(\boldsymbol{\theta})$ to approxmiate the posterior $p(\boldsymbol{\theta}|\mathcal{U})$. The approximation process is as follows.
@@
where $\mathbb{Q}$ is the distribution searching space and $Z_{t}$ is the intractable normalizing hyperparameter. Thus, referring to Appendix B, the ELBO is computed by
${\text{ELBO}}=-\lambda\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}H(\mathcal{U}_{t}|\boldsymbol{\theta})-\text{KL}\left(q(\boldsymbol{\theta})||p_{t}(\boldsymbol{\theta})\right).$ (6)
Optimizing with Eq. (6) may still suffer from large error accumulation. The entropy term may result in overly confident predictions that are less calibrated, while the KL term may be affected by an unreliable prior. In the following section, we will discuss how to solve the problems when computing the two terms.
4 Adaptation and Inference in VCoTTA
4.1 Mean-Teacher architecture
In the above section, we introduce the VI in CTTA but challenges remain, i.e., the prior drift. To mitigate the challenge, we adopt a Mean-Teacher (MT) structure (Tarvainen & Valpola, 2017) in the Bayesian inference process. MT is initially proposed in semi-supervised and unsupervised learning, where the teacher model guides the unlabeled data, helping the model generalize and improve performance with the utilization of large-scale unlabeled data.
MT structure is composed of a student model and a teacher model, where the student model learns from the teacher and the teacher updates using Exponential Moving Average (EMA) (Hunter, 1986). In this work, the student is set to be the variational distribution $q(\boldsymbol{\theta})$, which is a Gaussian mean-field approximation for its simplicity. It is achieved by stacking the biases and weights of the network as follows.

### v2

@@ Abstract
Continual Test-Time Adaptation (CTTA) task investigates effective domain adaptation under the scenario of continuous domain shifts during testing time. Due to the utilization of solely unlabeled samples, there exists significant uncertainty in model updates, leading CTTA to encounter severe error accumulation issues. In this paper, we introduce VCoTTA, a variational Bayesian approach to measure uncertainties in CTTA. At the source stage, we transform a pretrained deterministic model into a Bayesian Neural Network (BNN) via a variational warm-up strategy, injecting uncertainties into the model. During the testing time, we employ a mean-teacher update strategy using variational inference for the student model and exponential moving average for the teacher model. Our novel approach updates the student model by combining priors from both the source and teacher models. The evidence lower bound is formulated as the cross-entropy between the student and teacher models, along with the Kullback-Leibler (KL) divergence of the prior mixture. Experimental results on three datasets demonstrate the method’s effectiveness in mitigating error accumulation within the CTTA framework.
Index Terms:
continual test-time adaptation, domain adaptation, variational inference, error accumulation.
I Introduction
When deploying a pre-trained model for real-world applications, environmental changes are common. Recently, test-time adaptation (TTA) [liu2023test, verma2025real] tasks have garnered scholarly attention and made significant progress, focusing on how pre-trained models can automatically adapt to domain changes during testing. However, traditional TTA is typically suited for single-domain changes, while the Continual Test Time Adaptation (CTTA) task extends this concept to long-term scenarios. CTTA [wang2022continual] aims to enable a model to accommodate a sequence of distinct distribution shifts during the testing time, making it applicable to various risk-sensitive applications in open environments, such as autonomous driving and medical imaging. However, real-world non-stationary test data exhibit high uncertainty in their temporal dynamics [huang2022extrapolative], presenting a major challenge named error accumulation [wang2022continual].
Figure 1: In CTTA task, a BNN model is first trained on a source dataset, and then is used to adapt to updated with unreliable priors, which may suffer from noisy update. In this paper, we use the Bayesian approach to measure the uncertainty and try to reduce the effect of unreliable priors, achieving better adaptation.
To solve the error accumulation issue, existing CTTA studies rely on methods that enforce prediction confidence, such as entropy minimization. The majority of the existing methods focus on improving the confidence of the source model during the testing phase. For example, some methods employ the mean-teacher architecture [tarvainen2017mean] to mitigate error accumulation, where the student learns to align with the teacher and the teacher updates via moving average with the student. As to the challenge of forgetting source knowledge, some methods adopt augmentation-averaged predictions [wang2022continual, brahma2023probabilistic, dobler2023robust, yang2023exploring] for the teacher model, strengthening the teacher’s confidence to reduce the influence from highly out-of-distribution samples. Some other methods, such as [dobler2023robust, chakrabarty2023sata], propose to adopt the contrastive loss to maintain the already learned semantic information. Moreover, some researchers believe that the source model is more reliable, thus they are designed to restore the source parameters [wang2022continual, brahma2023probabilistic]. Though the above methods keep the model from confusion of vague pseudo labels, they may suffer from overly confident predictions that are less calibrated. As a result, these methods cannot adaptively distinguish between reliable and unreliable samples, which leaves them prone to error accumulation and limits their ability to quantify risks during prediction.
@@ I Introduction
In CTTA, the reliable estimation of uncertainty becomes particularly crucial in the context of continual distribution shift [ovadia2019can]. It is meaningful to design a model capable of encoding the uncertainty associated with temporal dynamics and effectively handling distribution shifts. To address this problem, we refer to the Bayesian Inference (BI) [box2011bayesian], which retains a distribution over model parameters that indicates the plausibility of different settings given the observed data, and it has been witnessed as effective in traditional continual learning (CL) tasks [nguyen2018variational], which also perform in long-term processes. In Bayesian continual learning, the posterior in the last learning task is set to be the current prior, which will be multiplied by the current likelihood. This kind of prior transmission is designed to reduce catastrophic forgetting in continual learning [wickramasinghe2023continual]. However, this is not feasible in CTTA because unlabeled data may introduce unreliable priors. As shown in Fig. 1, an unreliable prior may lead to a poor posterior, which may then propagate errors to the next inference, leading to the issue of error accumulation. The objective of this paper is to devise a CTTA procedure that not only enhances predictive accuracy under distribution shifts but also provides reliable uncertainty estimates by using the Bayesian method.
@@
Thus, we delve into the utilization of BI framework to evaluate model uncertainty in CTTA, aiming to mitigate the impact of unreliable priors and reduce the error propagation. To approximate the intractable likelihood in BI, we adopt to use online Variational Inference (VI) [wang2011online, sato2001online], and accordingly name our method Variational Continual Test-Time Adaptation (VCoTTA). At the source stage, we first transform a pretrained deterministic model into a Bayesian Neural Network (BNN) by a variational warm-up strategy, where the local reparameterization trick [kingma2015variational] is used to inject uncertainties into the source model. During the testing phase, we employ a mean-teacher update strategy, where the student model is updated via VI and the teacher model is updated by the exponential moving average. Specifically, for the update of the student model, we propose to use a mixture of priors from both the source and teacher models, then the Evidence Lower BOund (ELBO) becomes the cross-entropy between the student and teachers plus the KL divergence of the prior mixture. We demonstrate the effectiveness of the proposed method on three benchmark datasets, where the results show that it can mitigate the error accumulation caused by unreliable priors in CTTA, leading to clear performance improvements.
@@
  1. (1)
    We propose VCoTTA, a general framework that formulates CTTA as online variational inference within a BNN. This formulation explicitly models uncertainty and mitigates the error accumulation that arises from unreliable priors under persistent distribution shifts.
  2. (2)
    We design a variational warm-up strategy that converts off-the-shelf pre-trained models into BNNs, enabling uncertainty injection while preserving prior knowledge. This allows effective adaptation without retraining from scratch.
  3. (3)
    We introduce an adaptive entropy-based prior mixture, which dynamically fuses the source prior and teacher prior to provide a more reliable updating mechanism. Unlike ensemble-based methods, our mixture weights are determined by uncertainty, ensuring robustness in continuously changing environments.
II Related Work
II-A Continual Test-Time Adaptation
@@
TTA enables the model to dynamically adjust to the characteristics of the test data, i.e. target domain, in a source-free and online manner [jain2011online, sun2020test, wang2020tent]. Recently, CTTA [wang2022continual] has been introduced to tackle TTA within a continuously changing target domain, involving long-term adaptation [kim2021domain]. This configuration often grapples with the challenge of error accumulation [tarvainen2017mean, wang2022continual]. Specifically, prolonged exposure to unsupervised loss from unlabeled test data during long-term adaptation may result in significant error accumulation. Additionally, as the model is intent on learning new knowledge, it is prone to forgetting source knowledge, which poses challenges when accurately classifying test samples similar to the source distribution.
Some methods employ the mean-teacher architecture [tarvainen2017mean] to mitigate error accumulation, where the student learns to align with the teacher and the teacher updates via moving average with the student. CoTTA [wang2022continual] used weight-averaged and augmentation-averaged predictions along with stochastic restoration of some neurons to the source pre-trained weights. ECoTTA [song2023ecotta] proposed a memory-efficient approach, which utilizes lightweight meta networks and a self-distilled regularization technique to minimize memory consumption. Some methods propose to adopt the contrastive loss to maintain the already learnt semantic information. RMT [dobler2023robust] utilized symmetric cross-entropy loss for mean teachers and contrastive learning to address error accumulation. SWA [yang2023exploring] introduced a framework that enhances pseudo-label learning through self-adaptive thresholding, soft-weighted contrastive learning, and soft weight alignment to address the challenges of noisy pseudo-labels and knowledge forgetting. Some methods believe that the source model is more reliable, thus they are designed to use the source parameters more effectively. Gan et al.[gan2023decorate] introduced a visual domain prompt approach, which learns lightweight image-level visual prompts to reformulate input data for adapting to changing target domains without fine-tuning the source model. PETAL [brahma2023probabilistic] proposed a probabilistic framework, which employs a student-teacher model, a regularizer based on the source model’s posterior, and a data-driven parameter restoration technique using Fisher information matrix. VIDA [liu2023vida] employed high-rank and low-rank embedding spaces to manage domain-specific and domain-shared knowledge, respectively, and utilizes a homeostatic knowledge allotment strategy to dynamically fuse the knowledge from these adapters for improved adaptation performance. ADMA [liu2024continual] employed distribution-aware masking and histograms of oriented gradients reconstruction to enhance target domain knowledge extraction. Though the above methods alleviate the confusion caused by vague pseudo labels, they often lead to over-confident predictions with poor calibration. This indicates that existing deterministic strategies lack mechanisms to reliably handle sample-wise uncertainty, leaving room for approaches that explicitly incorporate uncertainty estimation.
II-B Bayesian Neural Network
@@
Bayesian framework is natural to incorporate past knowledge and sequentially update the belief with new data [zhao2022deep]. The bulk of work on Bayesian deep learning has focused on scalable approximate inference methods. These methods include stochastic VI [hernandez2015probabilistic, louizos2017multiplicative], dropout [gal2016dropout, kingma2015variational] and Laplace approximation [ritter2018scalable, friston2007variational] etc., and leveraging the stochastic gradient descent (SGD) trajectory, either for a deterministic approximation or sampling. In a BNN, we specify a prior $p(\boldsymbol{\theta})$ over the neural network parameters, and compute the posterior distribution over parameters conditioned on training data, $p(\boldsymbol{\theta}|\mathcal{D})\propto p(\boldsymbol{\theta})p(\mathcal{D}|\boldsymbol{\theta})$. This procedure should give considerable advantages for reasoning about predictive uncertainty, which is especially relevant in the small-data setting.
@@
Crucially, when performing BI, we need to choose a prior distribution that accurately reflects the prior beliefs about the model parameters before seeing any data [gelman1995bayesian, fortuin2021bayesian]. In conventional static machine learning, the most common choice for the prior distribution over the BNN weights is the simplest one: the isotropic Gaussian distribution. However, this choice has been proved indeed suboptimal for BNNs [fortuin2021bayesian]. Recently, some studies estimate uncertainty in continual learning within a BNN framework, such as [nguyen2018variational, ebrahimi2019uncertainty, farquhar2019unifying, kurle2019continual]. They set the current prior to the previous posterior to mitigate catastrophic forgetting. However, the prior transmission is not reliable in the unsupervised CTTA task. Any prior mistakes will be enlarged by adaptation progress, manifesting error accumulation. To solve the unreliable prior problem, this paper proposes a prior mixture method based on VI.
III Variational Inference in CTTA
III-A BI in traditional CL and in CTTA
@@
We start from the supervised BI in typical continual learning, where the model aims to learn multiple classification tasks in sequence. Let $\mathcal{D}=\{(x_{n},y_{n})\}_{n=1}^{N}$ be the training set, where $x_{n}$ and $y_{n}$ denotes the training sample and the corresponding class label. The task $t$ is to learn a direct posterior approximation over the model parameter $\boldsymbol{\theta}$ as follows.
@@
where $p(\boldsymbol{\theta}|\mathcal{D}_{1:t})$ denotes the posterior of sequential tasks on the learned parameter and $p(\mathcal{D}_{t}|\boldsymbol{\theta})$ is the likelihood of the current task. The current prior $p_{t}(\boldsymbol{\theta})$ is regarded as the given knowledge. [nguyen2018variational] proposes that this current prior can be the posterior learned in the last task, i.e., $p_{t}(\boldsymbol{\theta})=p(\boldsymbol{\theta}|\mathcal{D}_{1:t-1})$, where the inference becomes
@@
In contrast to continual learning [dai2025prompt], CTTA faces a sequence of learning tasks in test time without any label information, requiring the model to adapt to each novel domain sequentially. In this case, we assume that each domain is i.i.d. and the classes are separable following many unsupervised studies [miller1996mixture, van2020survey, caron2018deep]. We use $\mathcal{U}=\{x_{n}\}_{n=1}^{N}$ to represent the unlabeled test dataset. The CTTA model is first trained on a source dataset $\mathcal{D}_{0}$, and then adapted to unlabeled test domains starting from $\mathcal{U}_{1}$. For the $t$-th adaptation, we have
@@
Similarly, we can set the last posterior to be the current prior, i.e., $p_{t}(\boldsymbol{\theta})=p(\boldsymbol{\theta}|\mathcal{U}_{1:t-1}\cup\mathcal{D}_{0})$ and $p_{1}(\boldsymbol{\theta})=p(\boldsymbol{\theta}|\mathcal{D}_{0})$. However, employing BI for adaptation on unlabeled testing data can result in untrustworthy posterior estimates. Therefore, during subsequent adaptation, the untrustworthy posterior automatically transforms into unreliable priors, leading to error accumulation. In other words, an unreliable prior $p_{t}(\boldsymbol{\theta})$ will make the current posterior even less trustworthy. Moreover, the joint likelihood $p(\mathcal{U}_{t}|\boldsymbol{\theta})$ for $t>0$ is intractable on unlabeled data. The BI in CL and in CTTA can be compared in Fig. 2. For CL, BI is conducted by the posterior propagation, that is, the prior of the next task is equal to the current posterior. This is feasible in supervised CL, where the data label is provided. For CTTA, the posterior is not trustworthy using only pseudo labels to adapt to a new domain. Thus, propagating the untrustworthy posterior to the next stage would make it unreliable prior, which will result in error accumulation.
(a) BI in continual learning
(b) BI in CTTA (Our method)
Figure 2: BI in continual learning versus CTTA. We find the traditional prior transmission is infeasible in CTTA because of the unreliable prior from unlabeled data. In our method, we place CTTA in a mean-teacher structure, and design BI in CTTA using a mixture of teacher prior and source prior. The next teacher prior is updated by the exponential moving average.
III-B VI in CTTA
To make the BI feasible in CTTA task, in this paper, we transform the question to an easy-to-compute form. Referring to [grandvalet2004semi], the unsupervised inference can be transformed into
@@ III-B VI in CTTA
where $H$ denotes the conditional entropy and $\lambda$ is a scalar hyperparameter to weigh the entropy term. This simple form reveals that the prior belief about the conditional entropy of labels is given by the inputs. The observation of the input $\mathcal{U}$ provides information on the drift of the input distribution, which can be used to update the belief over the learned parameters $\boldsymbol{\theta}$ through Eq. (4). Consequently, this allows the utilization of unlabeled data for BI. More detailed derivations can be seen in Section V-A.
In a BNN, the posterior distribution is often intractable and some approximation methods are required, even when calculating the initial posterior. In this paper, we leverage online VI, as it typically outperforms the other methods for complex models in the static setting [bui2016deep]. VI defines a variational distribution $q(\boldsymbol{\theta})$ to approxmiate the posterior $p(\boldsymbol{\theta}|\mathcal{U})$. The approximation process is as follows.
@@
where $\mathbb{Q}$ is the distribution searching space and $Z_{t}$ is the intractable normalizing hyperparameter. Specifically, the KL divergence is expanded as
$\mathrm{KL}\left[q(\boldsymbol{\theta})\parallel\cfrac{1}{Z_{t}}p_{t}(\boldsymbol{\theta})e^{-\lambda H(\mathcal{U}_{t}|\boldsymbol{\theta})}\right]$ (6)
$=$ $-\int_{\boldsymbol{\theta}}q(\boldsymbol{\theta})\log\cfrac{\cfrac{1}{Z_{t}}p_{t}(\boldsymbol{\theta})e^{-\lambda H(\mathcal{U}_{t}|\boldsymbol{\theta})}}{q(\boldsymbol{\theta})}d\boldsymbol{\theta}$
$=$ $\int_{\boldsymbol{\theta}}q(\boldsymbol{\theta})\log Z_{t}d\boldsymbol{\theta}+\lambda\int_{\boldsymbol{\theta}}q(\boldsymbol{\theta})H(\mathcal{U}_{t}|\boldsymbol{\theta})d\boldsymbol{\theta}$
$-\int_{\boldsymbol{\theta}}q(\boldsymbol{\theta})\log\cfrac{p_{t}(\boldsymbol{\theta})}{q(\boldsymbol{\theta})}d\boldsymbol{\theta}$
where the first constant term can be reduced in the optimization. Thus, the ELBO can be computed by
${\text{ELBO}}=-\underbrace{\lambda\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}H(\mathcal{U}_{t}|\boldsymbol{\theta})}_{{\color[rgb]{0,0,0}\text{Entropy term}}}-\underbrace{\text{KL}\left(q(\boldsymbol{\theta})||p_{t}(\boldsymbol{\theta})\right)}_{{\color[rgb]{0,0,0}\text{KL term}}}.$ (7)
Optimizing with Eq. (7) makes the model adapt to domain shift. While VI offers a good framework for measuring uncertainty in CTTA, it is noteworthy that VI does not directly address the issue of unreliable priors. The error accumulation remains a significant concern.
Despite this, the form of the ELBO in variational inference offers a pathway for mitigating the impact of unreliable priors. In Eq. (7), the entropy term may result in overly confident predictions that are less calibrated, while the KL term may be directly affected by an unreliable prior. In the following section, we will discuss how to solve the problems when computing the two terms.
IV Adaptation and Inference in VCoTTA
IV-A Entropy term: Integrating Mean Teacher into VI
In the above section, we show that unreliable priors are a key challenge in VI-based CTTA. To address this, our solution is to design an uncertainty-aware prior mixture, where uncertainty determines how much the model should rely on the source prior or the teacher prior. To implement this mechanism in practice, we adopt the standard Mean Teacher (MT) framework [tarvainen2017mean] as a backbone, following prior works in CTTA [wang2022continual]. It is important to note that MT itself is not our contribution, but a commonly used mechanism that provides the teacher distribution for our Bayesian formulation. MT is initially proposed in semi-supervised and unsupervised learning, where the teacher model guides the unlabeled data, helping the model generalize and improve performance with the utilization of large-scale unlabeled data. Because of these advantages, MT is used in some CTTA methods such as [wang2022continual], thus we also build our method on it.
MT structure is composed of a student model and a teacher model, where the student model learns from the teacher and the teacher updates using Exponential Moving Average (EMA) [hunter1986exponentially]. In VI, the student is set to be the variational distribution $q(\boldsymbol{\theta})$, which is a Gaussian mean-field approximation for its simplicity. It is achieved by stacking the biases and weights of the network as follows.